OpenStudy (anonymous):

one more question....x(x-3)=x squared+6x+7

OpenStudy (anonymous):

expand the left side then you arrange everything to one side and you get yourself a quadratic equation should be easy to solve then

OpenStudy (anonymous):

u just comfused me

OpenStudy (anonymous):

expanding is multiplying what is in the parenthesis with the number outside.

OpenStudy (anonymous):

confused**

OpenStudy (anonymous):

so you're going to do x times x, which is x squared. then x times -3 to get -3x. you then solve for x

OpenStudy (anonymous):

x(x-3)=x^2-3x now x^2-3x=x^2+6x+7 this will simplify to: 3x+7=0... easy to solve now?

OpenStudy (anonymous):

right so the out come would be x*x-x*3=x*x+6x+7??

OpenStudy (anonymous):

yes :]

OpenStudy (anonymous):

or x^2-3x=x^2+6x+7 like josiahh said.

OpenStudy (anonymous):

okkk thanks yall!

OpenStudy (anonymous):

show some medal love!

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

clicking best response :P it gets peoples grades up!

OpenStudy (anonymous):

oh okay